# Weaving Geodesics and New Phenomena in Horocyclic Dynamics

### v2

@@ 2. Minimal Orbit Closures
The main part of the claim is the $N$-minimality. Our strategy will be to show that all accumulation points of $Nx_{0}$ are $A$-proximal to $x_{0}$ therefore by 2.3 have dense $N$-orbits in $F$.
Let $y\in F$ be an accumulation point of $Nx_{0}$ and let $0<\varepsilon<1$ be arbitrary. By Facts 1.3 and 1.4 we know $\mathscr{S}_{+}(A_{+}y)<\infty$ and that for any $\varepsilon>0$ a subray of $A_{+}y$ is contained in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$. Let $T_{1}>0$ be such that $\mathscr{S}_{+}(A_{[T_{1},\infty)}y)<\varepsilon$ and $A_{[T_{1},\infty)}y$ is contained in the $\varepsilon$-neighborhood of $\mathsf{T}^{1}_{+}\lambda$.
[…]
Now let $nx_{0}$ be sufficiently close to $y$ so as to satisfy $d(a_{T_{2}}nx_{0},a_{T_{2}}y)<\varepsilon$, for $T_{2}$ above, and also $d(nx_{0},y)<1$. Set $w=a_{T_{2}}nx_{0}$. Recall that the function $\beta$ is $N$-invariant and hence $\beta(nx_{0})=\beta(x_{0})=0$ implying that $\mathscr{S}_{+}(A_{+}nx_{0})=\tau(nx_{0})$. Since $\tau$ is 1-Lipschitz we have $\tau(nx_{0})<\tau(y)+1$ and hence
$\mathscr{S}_{+}(A_{+}w)\leq\mathscr{S}_{+}(A_{+}nx_{0})<\tau(y)+1=\rho.$
By our choice of $T_{2}$ we conclude that $A_{+}w$ is weaving with pattern $W\geq k_{0}$. That is, $A_{+}w$ is asymptotic to some $\eta_{W,\pm}$ and moreover $d(w,\eta_{W,\pm})<2\varepsilon$. This implies that
(2.3) $\mathscr{S}_{+}(A_{+}w)<2\varepsilon+\mathscr{S}_{+}(\eta_{W,\pm})<3\varepsilon.$
But recall that $d(a_{t}nx_{0},a_{t}x_{0})\to 0$ as $t\to\infty$ and hence $d(a_{t-T_{2}}w,a_{t}x_{0})\to 0$. Joining this fact with (2.3) we deduce that $|\tau(w)-T_{2}|<3\varepsilon$.
Since $d(w,a_{T_{2}}y)<\varepsilon$ we have $|\tau(a_{T_{2}}y)-T_{2}|<4\varepsilon$. On the other hand, by our choice of $T_{2}>T_{1}$ we had
$\mathscr{S}_{+}(A_{[T_{2},\infty)}y)<\varepsilon.$
We thus conclude $|\beta(a_{T_{2}}y)-T_{2}|<5\varepsilon$, or in other words, $|\beta(y)|<5\varepsilon$ which by Lemma 2.4 implies $y$ is $A$-proximal to $a_{b}x_{0}$ for some $|b|<5\varepsilon$ (in fact, necessarily $0\leq b$). Since $\varepsilon$ was arbitrary we conclude the claim. ∎
@@
Note that minimality together with (1.2) implies that $\beta$ is constant (zero) on all of $\overline{Nx_{0}}$. This also follows directly from our proof. Moreover, the fact that $\beta(a_{t}z)=\beta(z)+t$ for all $t\in\mathbb{R}$ implies that $a_{t}y\notin\overline{Nx_{0}}$ for all $y\in\overline{Nx_{0}}$ and $t\neq 0$.
@@ 2.1. Invariant Measures
Proof of 2.7.
The existence of an $N$-invariant locally finite measure supported on $\overline{Nx_{0}}$ follows from [KMR13, Lemma 2.2] and the superamenability of the group $\mathbb{R}$ [Ros74]. See also [Ami22]. For completeness and accessibility, we provide a short, more hands-on, construction of such measures in A.1 in the appendix.
These statements ensure the existence of an $N$-invariant, conservative and locally-finite measure supported on $\overline{Nx_{0}}$. Let $\mu$ denote a locally-finite ergodic component of said measure. By Ratner’s classification of finite $N$-invariant measures [Rat92] (and the fact that $\overline{Nx_{0}}$ is not a homogeneous subspace of $\mathsf{T}^{1}\Sigma_{s}$) we conclude $\mu$ is necessarily infinite.
@@
We note that the notion of a “trivial” measure in this paper is strictly broader than the one used by Sarig in [Sar10], who referred only to measures corresponding to periodic horocycles or orbits based outside the limit set. The measures constructed in [Sar10, Theorem 3] are trivial in the sense that they are supported on single proper horocycle orbits, which was known by the author.
The measures constructed in [Sar10, Theorem 3] are ergodic components of Lebesgue measure on certain surfaces given by Fuchsian groups of the first kind with critical exponent<1/2<\!1/2, see [Pat79]. In [Pat77, Equation (18)], Patterson shows that Fuchsian groups $\Gamma$ having critical exponent smaller than one-half satisfy, in the Poincaré disk model,
@@
Recall that in the disk model one has $|\gamma^{\prime}(\xi)|=e^{-\beta_{\xi}(\gamma^{-1}.0,0)}$ where $\beta$ is the Busemann cocycle, see e.g. [Sar19, Appendix 1]. Therefore, we have that for Leb-a.e. $\xi\in\partial\mathbb{H}^{2}$, any horocycle tangent to $\xi$ has only finitely many points in $\Gamma.0$ within a bounded distance of it. This in turn implies that such horocycles have no accumulation points in the quotient surface.
@@ 3. Fractional Hausdorff Dimension
See [FLM24, §2.2] and [FLM23, §7] for another account of such sets (under different notations).
The main theorem in this section is the following:
@@
The vast flexibility of our construction allows us to exhibit any compact subset of $(0,\infty)$ as $E$, that is, we have:
@@
For any compact [^2] Our definition of loom surfaces is not the most general imaginable. In particular, the assumption that the $h_{k}$ are bounded is merely imposed for simplicity. One can construct surfaces with an unbounded sequence of $h_{k}$ under which additional conditions on $|s_{k+1}-s_{k}|\gg 0$ would ensure all the claims of this paper will hold. set $E\subset(0,\infty)$ there exists a distal loom surface $\Sigma_{s}$ in which $E=\mathrm{accum}\left(\mathscr{S}_{+}(\eta_{k}^{\pm})\right)_{k\in\mathbb{N}}$.
@@
Given $E$, choose a finite or countable dense subset $\{e_{1},e_{2},$...\} of $E$. By solving for $h$ in 1.7 we can construct $\Sigma_{s}$ where every $e_{j}$ appears infinitely many times as the slack of different crossings. The proposition follows. ∎
@@ 3.1.1. Fixed Fractional Hausdorff Dimension
For any $\alpha\in[0,1]$ there exists a compact set $E^{\prime}$ with $\dim_{\mathrm{H}}mE^{\prime}=\alpha$ for all $m\geq 1$, see [SS10, Kör08]. Since $E^{\prime}$ may contain 0, and the accumulations of slacks of geodesics $\eta_{k}^{\pm}$ are all strictly positive in a distal loom surface, we take $E=1+E^{\prime}$. By Proposition 3.5, there is a distal loom surface with $E=\mathrm{accum}\left(\mathscr{S}_{+}(\eta_{k}^{\pm})\right)_{k\in\mathbb{N}}$. Since $\dim_{\mathrm{H}}mE=\dim_{\mathrm{H}}mE^{\prime}$, Theorem 3.3 gives the following:
@@ 3.1.2. Locally Varying Dimensions
In fact, in [SS10] the authors construct compact sets $E^{\prime}\subset[0,1]$ satisfying:
@@
are arbitrary non-decreasing sequences (where $\{\beta_{m}\}$ and $\{\gamma_{m}\}$ are required to satisfy some mild growth constraints). See also [Kör08].
@@ 3.1.3. Discrete Sub-Invariance
see [FLM23, §7]. By constructing $\Sigma_{s}$ with $E=\{T_{0}\}$ for some $T_{0}>0$ we obtain a horocyclic orbit closure with a discrete semigroup of sub-invariance. This is in contrast with orbit closures in $\mathbb{Z}$-covers of compact surfaces, see [FLM23, Prop. 7.20].
@@
Let $t\in 2kE$ with $t=\sum_{j=1}^{2k}e_{j}$ where $e_{j}\in E$. Let $\varepsilon>0$ and let $S>0$ be the constant in 1.9 corresponding to $\varepsilon$ and $m=2k$. Let $(s_{k})$ be the sequence of horizontal positions of boundary components in the definition of $\Sigma_{s}$. Let $T>0$ be large enough that $|s_{k}-s_{l}|>S$ holds for all distinct $s_{k},s_{l}>T$. By the definition of $E$ there exists for each $e_{j}$ a crossing $\eta^{\pm}_{k_{j}}$ with $|\mathscr{S}_{+}(\eta_{k_{j}}^{\pm})-e_{j}|<\varepsilon/2k$ and with $s_{k_{j}}>T$. Hence the weaving geodesic $\eta_{W}^{+}$ with weaving pattern $\{k_{1},...k_{2k}\}$ satisfies the conditions of 1.9 implying:
@@ Appendix A Construction of Invariant Measure
We now focus on the particular setting of this paper. Let $\Sigma_{s}$ be a loom surface satisfying the summability condition and let $Y=\overline{Nx_{0}}$ be $N$-minimal as discussed above. We will deduce the existence of a locally finite invariant measure supported on $Y$ from the existence of a certain section to $Y$, that is, a subset $\Psi\subseteq\mathsf{T}^{1}\Sigma_{s}$ for which $\left\{n\in N:nz\in\Psi\right\}$ is a discrete (possibly empty) subset of $N$ for all $z\in\mathsf{T}^{1}\Sigma_{s}$.
@@
There exists a section $\Psi$ for the $N$-flow satisfying:
@@
This lemma together with A.1 concludes the proof of the existence of the desired measure.
@@
Recall that $NAU$ corresponds to the open Bruhat cell in $\operatorname{PSL}_{2}(\mathbb{R})$, which means in particular that the multiplication map $N\times A\times U\to NAU$ is a diffeomorphism (see e.g. [Kna02, Lemma 6.44] and §2.3 in [FLM23] for more details). By our choice of constants and the fact that the parameterizations of $U$, NN and $A$ are of unit speed, we conclude that $N_{(-\delta/4,\delta/4)}\Psi\subset B_{\delta}^{\mathsf{T}^{1}\Sigma_{s}}(x_{0})$ hence showing the multiplication map $N_{(-\delta/4,\delta/4)}\times A_{(-\delta/4,\delta/4)}\times U_{(c,d)}\to N_{(-\delta/4,\delta/4)}\Psi$ is a homeomorphism.

### v3

@@ 2. Minimal Orbit Closures
The main part of the claim is the $N$-minimality. As promised, our strategy will be to show that all accumulation points of $Nx_{0}$ are $A$-proximal to $x_{0}$ and therefore, by 2.3, have dense $N$-orbits in $F$. By 2.5, it suffices to prove that $\beta(y)=0$ for all $y\in F$, which by definition amounts to showing:
$\mathscr{S}_{+}(A_{+}y)=\tau(y).$
Let $y\in F$ and let $0<\varepsilon<1$ be smaller than half the minimal injectivity radius in $\Sigma_{s}$. By 1.6 and Lemma 2.6 there exists $k_{0}$ large enough so that any geodesic ray $A_{+}w$ beginning at $\tau(w)\geq s_{k_{0}}$ with $\mathscr{S}_{+}(A_{+}w)\leq\rho=\tau(y)+1$ is weaving with pattern $W\geq k_{0}$ for which
(2.4) $\mathscr{S}_{+}(\eta_{W,\pm})<\varepsilon.$
By increasing $k_{0}$ if necessary and using Lemma 2.4 we may ensure $\delta_{k}<\varepsilon$ for all $k\geq k_{0}$.
Now fix $T\geq 0$ for which $d_{\mathsf{T}^{1}\Sigma_{s}}(a_{T}y,a_{s_{k}}x_{0})<\varepsilon$ with $\tau(a_{T}y)=s_{k}>s_{k_{0}}+1$ for some $k>k_{0}$ as ensured by Lemma 2.4. By increasing $T$ if necessary we may further require that $\mathscr{S}_{+}(A_{[T,\infty)}y)<\varepsilon$, as this expression decreases to 0 in $T$.
Let $nx_{0}\in Nx_{0}$ be so close to $y$ so as to satisfy
(2.5) $d_{\mathsf{T}^{1}\Sigma_{s}}(nx_{0},y)<\varepsilon\quad\text{and}\quad d_{\mathsf{T}^{1}\Sigma_{s}}(a_{T}nx_{0},a_{T}y)<\varepsilon.$
Denote $w=a_{T}nx_{0}$. The function $\beta$ being $N$-invariant implies $\beta(nx_{0})=\beta(x_{0})=0$ and hence
(2.6) $\mathscr{S}_{+}(A_{+}nx_{0})=\tau(nx_{0}).$
We thus have
$\mathscr{S}_{+}(A_{+}w)$ $\leq\mathscr{S}_{+}(A_{+}nx_{0})=\tau(nx_{0})\leq$
$\leq\tau(y)+d_{\mathsf{T}^{1}\Sigma_{s}}(nx_{0},y)\leq\tau(y)+1=\rho,$
by the 1-Lipschitz property of $\tau$. Since $\tau(w)>s_{k}-1>s_{k_{0}}$ we are ensured by Lemma 2.6 that $A_{+}w$ is weaving of pattern $W\geq k_{0}$ satisfying (2.4). Let $\eta_{W,\sigma}$ correspond to $A_{+}w$ as in Lemma 1.7. By Lemma 2.4 and our choice of $a_{T}y$ we know that
$d_{\Sigma_{s}}(p(w),\eta_{W,\sigma})\leq d_{\Sigma_{s}}(p(w),p(a_{T}y))+d_{\Sigma_{s}}(p(a_{T}y),\eta_{W,\sigma})<\varepsilon+\delta_{k}<2\varepsilon.$
Let $\alpha$ denote the shortest path connecting $p(w)$ to $\eta_{W,\sigma}$, then $\mathrm{length}(\alpha)<2\varepsilon$ is smaller than the injectivity radius around $p(w)$ implying that $A_{+}w$ is given by pulling tight $\alpha\ast\eta_{W,\sigma}$. We thus conclude
$\mathscr{S}_{+}(A_{+}w)$ $\leq 2\mathrm{length}(\alpha)+\mathscr{S}_{+}(\eta_{W,\sigma})<4\varepsilon+\varepsilon<5\varepsilon,$
where we made use of the fact that the slack of a path is always bounded above by twice its length, since $\tau$ is $1$-Lipschitz.
On the other hand, by (2.5) we have
$|\mathscr{S}_{+}(A_{[0,T]}y)-\mathscr{S}_{+}(A_{[0,T]}nx_{0})|\leq|\tau(y)-\tau(nx_{0})|+|\tau(a_{T}y)-\tau(a_{T}nx_{0})|<2\varepsilon.$
We therefore obtain
$|\mathscr{S}_{+}(A_{+}y)-\mathscr{S}_{+}(A_{+}nx_{0})|$ $\leq 2\varepsilon+\mathscr{S}_{+}(A_{[T,\infty)}y)+\mathscr{S}_{+}(A_{[T,\infty)}nx_{0})=$
$=2\varepsilon+\varepsilon+\mathscr{S}_{+}(A_{+}w)<8\varepsilon.$
By (2.6) we thus have
$|\mathscr{S}_{+}(A_{+}y)-\tau(y)|\leq|\mathscr{S}_{+}(A_{+}y)-\mathscr{S}_{+}(A_{+}nx_{0})|+|\tau(nx_{0})-\tau(y)|<9\varepsilon.$
Since $\varepsilon$ was arbitrary we conclude $\mathscr{S}_{+}(A_{+}y)=\tau(y)$, as claimed. ∎
@@
We have shown that $\beta$ is constant (zero) on all of $\overline{Nx_{0}}$. This also follows from the minimality of $\overline{Nx_{0}}$ together with (1.3). Moreover, the fact that $\beta(a_{t}z)=\beta(z)+t$ for all $t\in\mathbb{R}$ implies that $a_{t}y\notin\overline{Nx_{0}}$ for all $y\in\overline{Nx_{0}}$ and $t\neq 0$.
@@ 2.1. Invariant Measures
Proof of 2.8.
The existence of an $N$-invariant locally finite measure supported on $\overline{Nx_{0}}$ follows from [15, Lemma 2.2] and the superamenability of the group $\mathbb{R}$ [31]. See also [1]. For completeness and accessibility, we provide a short, more hands-on, construction of such measures in A.1 in the appendix.
These statements ensure the existence of an $N$-invariant, conservative and locally-finite measure supported on $\overline{Nx_{0}}$. Let $\mu$ denote a locally-finite ergodic component of said measure. By Ratner’s classification of finite $N$-invariant measures [29] (and the fact that $\overline{Nx_{0}}$ is not a homogeneous subspace of $\mathsf{T}^{1}\Sigma_{s}$) we conclude $\mu$ is necessarily infinite.
@@
We note that the notion of a “trivial” measure in this paper is strictly broader than the one used by Sarig in [33], who referred only to measures corresponding to periodic horocycles or orbits based outside the limit set. The measures constructed in [33, Theorem 3] are trivial in the sense that they are supported on single proper horocycle orbits, which was known by the author.
The measures constructed in [33, Theorem 3] are ergodic components of Lebesgue measure on certain surfaces given by Fuchsian groups of the first kind with critical exponent<1/2<\!1/2, see [27]. In [26, Equation (18)], Patterson shows that Fuchsian groups $\Gamma$ having critical exponent smaller than one-half satisfy, in the Poincaré disk model,
@@
Recall that in the disk model one has $|\gamma^{\prime}(\xi)|=e^{-\beta_{\xi}(\gamma^{-1}.0,0)}$ where $\beta$ is the Busemann cocycle, see e.g. [34, Appendix 1]. Therefore, we have that for Leb-a.e. $\xi\in\partial\mathbb{H}^{2}$, any horocycle tangent to $\xi$ has only finitely many points in $\Gamma.0$ within a bounded distance of it. This in turn implies that such horocycles have no accumulation points in the quotient surface.
@@ 3. Fractional Hausdorff Dimension
See [11, §2.2] and [10, §7] for another account of such sets (under different notations).
The main theorem of this section, proven in §3.2, is the following:
@@
The vast flexibility of our construction allows us to exhibit any compact [^2] Our definition of loom surfaces is not the most general imaginable. In particular, the assumption that the $h_{k}$ are uniformly bounded away from $\frac{\pi}{2}$ is merely imposed for simplicity. One can construct surfaces with an unbounded sequence of $h_{k}$ under which additional conditions on $|s_{k+1}-s_{k}|\gg 0$ would ensure all claims in this paper will hold. subset of $(0,\infty)$ as $E$, that is, we have:
@@
For any compact set $E\subset(0,\infty)$ there exists a distal loom surface $\Sigma_{s}$ in which $E=\mathrm{accum}\left(\mathscr{S}_{+}(\eta_{k}^{\pm})\right)_{k\in\mathbb{N}}$.
@@
Given $E$, choose a finite or countable dense subset $\{e_{1},e_{2},$...\} of $E$. By solving for $h$ in 1.8 we can construct $\Sigma_{s}$ where every $e_{j}$ appears infinitely many times as the slack of different crossings. The proposition follows. ∎
@@ 3.1.1. Fixed Fractional Hausdorff Dimension
For any $\alpha\in[0,1]$ there exists a compact set $E^{\prime}$ with $\dim_{\mathrm{H}}mE^{\prime}=\alpha$ for all $m\geq 1$, see [35, 17]. Since $E^{\prime}$ may contain 0, and the accumulations of slacks of geodesics $\eta_{k}^{\pm}$ are all strictly positive in a distal loom surface, we take $E=1+E^{\prime}$. By Proposition 3.5, there is a distal loom surface with $E=\mathrm{accum}\left(\mathscr{S}_{+}(\eta_{k}^{\pm})\right)_{k\in\mathbb{N}}$. Since $\dim_{\mathrm{H}}mE=\dim_{\mathrm{H}}mE^{\prime}$, Theorem 3.3 gives the following:
@@ 3.1.2. Locally Varying Dimensions
In fact, in [35] the authors construct compact sets $E^{\prime}\subset[0,1]$ satisfying:
@@
are arbitrary non-decreasing sequences (where $\{\beta_{m}\}$ and $\{\gamma_{m}\}$ are required to satisfy some mild growth constraints). See also [17].
@@ 3.1.3. Discrete Sub-Invariance
see [10, §7]. By constructing $\Sigma_{s}$ with $E=\{T_{0}\}$ for some $T_{0}>0$ we obtain a horocyclic orbit closure with a discrete semigroup of sub-invariance. This is in contrast with orbit closures in $\mathbb{Z}$-covers of compact surfaces, see [10, Prop. 7.20].
@@
Let $t\in 2kE$ with $t=\sum_{j=1}^{2k}e_{j}$ where $e_{j}\in E$. Let $\varepsilon>0$ and let $S>0$ be the constant in 1.10 corresponding to $\varepsilon$ and $m=2k$. Let $(s_{k})$ be the sequence of horizontal positions of boundary components in the definition of $\Sigma_{s}$. Let $T>0$ be large enough that $|s_{k}-s_{l}|>S$ holds for all distinct $s_{k},s_{l}>T$. By the definition of $E$ there exists for each $e_{j}$ a crossing $\eta^{\pm}_{k_{j}}$ with $|\mathscr{S}_{+}(\eta_{k_{j}}^{\pm})-e_{j}|<\varepsilon/2k$ and with $s_{k_{j}}>T$. Hence the weaving geodesic $\eta_{W}^{+}$ with weaving pattern $\{k_{1},...k_{2k}\}$ satisfies the conditions of 1.10 implying:
@@ Appendix A Construction of Invariant Measure
We now focus on the particular setting of this paper. Let $\Sigma_{s}$ be a loom surface satisfying the summability condition and let $Y=\overline{Nx_{0}}$ be $N$-minimal as discussed above.
@@
There exists a section $\Psi\subseteq\mathsf{T}^{1}\Sigma_{s}$ for the $N$-flow satisfying:
@@
This lemma together with A.1 concludes the proof of the existence of an $N$-invariant locally finite conservative measure, as desired.
@@
Recall that $NAU$ corresponds to the open Bruhat cell in $\operatorname{PSL}_{2}(\mathbb{R})$, which means in particular that the multiplication map $N\times A\times U\to NAU$ is a diffeomorphism (see e.g. [16, Lemma 6.44] and §2.3 in [10] for more details). By our choice of constants and the fact that the parameterizations of $U$, NN and $A$ are of unit speed, we conclude that $N_{(-\delta/4,\delta/4)}\Psi\subset B_{\delta}^{\mathsf{T}^{1}\Sigma_{s}}(x_{0})$ hence showing the multiplication map $N_{(-\delta/4,\delta/4)}\times A_{(-\delta/4,\delta/4)}\times U_{(c,d)}\to N_{(-\delta/4,\delta/4)}\Psi$ is a homeomorphism.